OpenStudy (anonymous):

According to the manual, a battery in a cellular phone loses 2% of its charge each day assume the battery is 100% charged a) Write an equation to represent the percent charge, P, as a function of the number of days, D, since the battery was charged. b) Determine the number of days until the battery in only 50% charged. Please show full solutions Thanks

OpenStudy (anonymous):

a: p=100 - (2xD) B: it would take 25 days to be at 50%

OpenStudy (anonymous):

can you please show me full work because i don't understand the question, thanks

OpenStudy (anonymous):

I think it's an exponential equation. Note it's discharging continuously.

OpenStudy (anonymous):

well in the equation p = 100 - (2D) P is the battery left, and it equals 100% - 2% battery lost in each day. so if 2 days went by, it would be 96% battery left because 96 = 100 - (2x2)

OpenStudy (anonymous):

I think its more like \[C=e ^{-.0202t}\] Where t is the number of days and C is the percent of charge expresses a decimal. So for a 50% charge, \[.5=e ^{-.0202t}\] t=34 days

OpenStudy (anonymous):

C is the percent of charge expressed as decimal.